OpenStudy (cutiekawdfi):

help please will fan and medal (((: find the real roots of each equation by graphing. 1. x^2 + 3 = x^3 - 5

OpenStudy (faiqraees):

make a graph of y=x^2 + 3 then make a graph of y=x^3 - 5 The intersection points are the roots

OpenStudy (cutiekawdfi):

i actually want to know how to do it..

OpenStudy (anonymous):

textbook lol

OpenStudy (jv0408):

graph it

OpenStudy (anonymous):

@faiqraees i think shes looking for the zeros of each equation

OpenStudy (cutiekawdfi):

youre not helping..

OpenStudy (anonymous):

i know...

OpenStudy (faiqraees):

make a graph of y=x^2 + 3 then make a graph of y=x^3 - 5 The intersection points are the roots Its the only way

OpenStudy (jv0408):

make a graph of y=x^2 + 3 then make a graph of y=x^3 - 5 The intersection points are the roots Its the only way

OpenStudy (faiqraees):

The zeroing technique will work by taking every element on the r.h.s and equate it to y and graph it but problem is it wont be fulfilling the question requirement

OpenStudy (anonymous):

@faiqraees, its not a system of equations

OpenStudy (anonymous):

top of the morning to you

OpenStudy (faiqraees):

Try it and you will have the answer 100 percent